OpenStudy (anonymous):

Whats the trick to solve for x with this equation? 0 = (2x-2x^3) E^(1-x^2) x = ???

OpenStudy (nincompoop):

what is this (2x-2x^3) E^(1-x^2)

OpenStudy (nincompoop):

use latex

OpenStudy (owlcoffee):

\[(2x-2x^3)e ^{(1-x^2)}\]

OpenStudy (anonymous):

its a derivative of f[x] = x^2 E^(-x^2​ + 1) apparently I have to examine the derivative and find the zeros, without a calculator, to plot a chart.. I can see just looking at the equations that 0,-1,1 will give me the zeros.. but I can't explain why that would be true..

OpenStudy (nincompoop):

why would you need a calculator?

OpenStudy (anonymous):

sorry, whats latex ?

OpenStudy (anonymous):

why a calc, coz I'm clueless of course..

OpenStudy (nincompoop):

you probably need to use log

OpenStudy (anonymous):

Yeah I was worried about that..

OpenStudy (anonymous):

To get rid of that E^ ?

OpenStudy (owlcoffee):

No, 1 and -1 are not zeros.

OpenStudy (nincompoop):

they are zeroes

OpenStudy (anonymous):

had me there for minute.. thought I sent the wrong equation

OpenStudy (nincompoop):

I am confused how you can "see" the values without solving it.

OpenStudy (owlcoffee):

He may have looked at the graphical representation.

OpenStudy (anonymous):

It just looks obvious to me.. I usually plug 1 and 0 in to get an idea of the function and if 1 zeros, -1 usually does too when squares are involved.

OpenStudy (nincompoop):

then use that as your mathematical solution "it is obvious"

OpenStudy (owlcoffee):

Nin, don't be so rude.

OpenStudy (nincompoop):

@UnkleRhaukus @UsukiDoll @rational

OpenStudy (owlcoffee):

if you want to find the zeroes of the function: \[(2x-2x^3)e ^{1-x^2}=0\] is clearly and application of the hankelian preoperty, wich means: \[2x-2x^3=0\] \[e ^{1-x^2}= 0\]

OpenStudy (anonymous):

thats a cool one.. first time I've heard of that.. but it makes sense yes.. if a*b = 0 then a or b = 0

OpenStudy (owlcoffee):

yes, now all you have to do is find the zeroes of those smaller and simpler parts of the function.

OpenStudy (anonymous):

silly question ... does that also apply for division?

OpenStudy (owlcoffee):

Not really, I think you mean: \[\frac{ a }{ b }=0 <=>a=0,b=0 \] Thing is that does not work, because the expression a/b has the followin condition: \[\frac{ a }{ b }=0 <=> a \in \mathbb{R}, b \neq 0\] because you can't have a zero in the denominator.

OpenStudy (amoodarya):

hint : \[e^{something}\neq0\]

OpenStudy (owlcoffee):

Do you mean that it does no have zeroes?

OpenStudy (anonymous):

ah amoodarya, thnx.. yea E^0 = 1

OpenStudy (nincompoop):

can you prove that \(e^{1-x^2} = 0 \)

OpenStudy (owlcoffee):

I can.

OpenStudy (anonymous):

(2x-2x^3) is all we need though right? because if that is 0 the whole equation is 0

OpenStudy (owlcoffee):

oh no, wait, log 0 is not possible.

OpenStudy (nincompoop):

I know it is not possible, but you said you can

OpenStudy (anonymous):

yeah, I tried to log both sides earlier.. when one side was 0, can you do that even temporarily ?

OpenStudy (owlcoffee):

Sorry, thought there would be a constant.

OpenStudy (nincompoop):

this is beyond my skill set

OpenStudy (anonymous):

so if the zero must be in the \[0 =(2x - 2x^3)\] then \[-2x = 2x - 2x^3 -2x\] \[-2x = - 2x^3 \] \[-2x/-2 = - 2x^3 /-2 \] \[x = x^3 \] \[x/x = x^3/x \] \[1 = x^2 \] \[1/x = x^2/x \] x = 1/x Does this look right?

OpenStudy (anonymous):

lol, it cant be

OpenStudy (owlcoffee):

Well, not really, just because : \[e ^{f(x)}=0\] Will always have positive images of x, because it is a lil' more complex than the expression: \[f:f(x)=e ^{x}\] therefore, the only part of the function that'll determine the zeroes will be: \[(2x-2x^3)\]

OpenStudy (nincompoop):

n0

OpenStudy (nincompoop):

I am interested to learn this

OpenStudy (owlcoffee):

To find the zeroes of the expression: \[2x-2x^3=0\] You take common factor 2x: \[2x(1-x^2)\] And again, by the hankelian propety: \[2x=0 \] \[1-x^2 =0\]

OpenStudy (anonymous):

I think the solution might also be here.. x = x^3 x = x x x

OpenStudy (nincompoop):

uh...

OpenStudy (anonymous):

lol, maybe not..

OpenStudy (nincompoop):

stop trolling

OpenStudy (usukidoll):

ummm.... is this solve for x or find the derivative ?

OpenStudy (anonymous):

solve for x

OpenStudy (usukidoll):

Alright... I'm going to use what @Owlcoffee suggested earlier.

OpenStudy (usukidoll):

\[(2x-2x^3)e ^{1-x^2}=0\] so we're going to solve for x... more like find all x's that satisfy the equation... this feels like find the equilibrium or something. \[(2x-2x^3)=0\] \[e ^{1-x^2}=0\]

OpenStudy (nincompoop):

good luck

OpenStudy (usukidoll):

for the first equation we have a 2x in common \[(2x-2x^3) \rightarrow 2x(1-x^2)\] factoring again yields \[2x(1+x)(1-x)=0\] just through analysis and looking at it x = 0,1,-1

OpenStudy (usukidoll):

\[e ^{1-x^2}=0\] if I take the ln on both sides because e^x and ln x are inverses of each other, that e should go away ... the problem is \[e^{\ln(1-x^2)}=\ln 0 \] 1-x^2 = ln 0 and ln 0 is undefined >:/

OpenStudy (usukidoll):

in fact ln (negative any number) is also undefined.

OpenStudy (nincompoop):

the easy part is always the best part, yes?

OpenStudy (owlcoffee):

Nin, what the hell? Usuki made a great job explaining it very very simple.

OpenStudy (usukidoll):

O____________________________________o

OpenStudy (anonymous):

would it not be okay though if \[E^(1-x^2) = 1\] because the zeros on the other term would satisfy what we need

OpenStudy (usukidoll):

do you mean \[e^{1-x^2} = 1\] ?

OpenStudy (anonymous):

yeah I think if x = 1,0,-1 that equation = 1 , and that's okay.

OpenStudy (anonymous):

or E

OpenStudy (nincompoop):

I still do not get E what is E

OpenStudy (anonymous):

the exponential constant = E

OpenStudy (nincompoop):

so it is not Euler's constant?

OpenStudy (usukidoll):

well let's supposed we are given that equation instead \[e^{1-x^2} = 1\] so taking the ln on both sides yields \[e^{\ln(1-x^2)} = \ln 1\]

OpenStudy (anonymous):

sorry, yes Eulers

OpenStudy (usukidoll):

? What does Euler's Method have to do with this? That's differential equations

OpenStudy (nincompoop):

you're dazzling as always, @UsukiDoll did you get your degree yet?

OpenStudy (usukidoll):

that question has nothing to do with what we are trying to solve. Please delete it.

OpenStudy (nincompoop):

:)

OpenStudy (owlcoffee):

If you know so much, Nin, why aren't you helping?

OpenStudy (usukidoll):

ok so e is the mathematical exponent.... it's also Euler's number... and it's default value is 2.71... but enough with that.. let's go back to what we're doing...this is going off topic and it's not helping hughfuve

OpenStudy (usukidoll):

so assuming we have \[e^{\ln(1-x^2)} = \ln 1\] since ln 1 -> 0 \[1-x^2=0\] \[(1+x)(1-x) = 0\] \[x=1,-1\]

OpenStudy (anonymous):

Is it a valid technique to solve for that first half of the equation... (2x - 2x^3) and then take the 3 zeros that you found.. and plug them into the second part of the equation. Or would that be frowned upon?

OpenStudy (usukidoll):

I'm trying to think if this is going to work... if I plug in x = 1 on the original equation... I will have \[e^{\ln 0} = 0\] I know ln 0 is undefined.. but \[e^x \] and \[\ln x \] are inverses of each other so I can use that to make the equation hold true. wait let me see that suggestion..

OpenStudy (usukidoll):

our zeros from the first equation x = 0,1,-1 x=1 is going to be a problem if I put it in the second equation

OpenStudy (usukidoll):

\[e ^{1-x^2}=0\] let x = 1 \[e ^{1-(1)^2}=0\] \[e ^{1-1}=0\] \[e ^{0}=0\] \[1 \neq 0\]

OpenStudy (usukidoll):

x = -1 will produce the same issue to because (-1)^2 - > (-1)(-1) = 1 Same thing will occur

OpenStudy (owlcoffee):

Thing is, if you plug the zeroes you found, say for instance, x=0: \[f(0)=(2(0)-2(0)^3)e ^{1-0^2}\] you'll end up having: \[f(0)=0.e^1\] which is just : \[f(0)=0\] And that'll happen for 1, and -1. it will make the whole function a "0", but it will not make e^(1-x^2) equal zero.

OpenStudy (usukidoll):

and we can't use x = 0 either you will end up with \[e^1 = 0 \rightarrow 2.71 \neq 0 \]

OpenStudy (anonymous):

but does E^(1-x^2) have to equal zero? if the full equation is \[0 == (2x-2x^3) (E^1-x^2)\] then even if E^(n) = 1 or E, is our equation not still a valid 0?

OpenStudy (anonymous):

oops on my equation.. its supposed to have everything after E in the exponent.

OpenStudy (usukidoll):

the zero's we found for x =0,-1,1 is not going to make \[e^{(1-x^2)}=0\] true for 2 reasons. you will either have \[e^0 \neq 0 \rightarrow 1 \neq 0\] or \[e \neq 0 \rightarrow 2.71 \neq 0 \]

OpenStudy (usukidoll):

but I think if you combine those two together... the left side of the equation the one with the 2x should make it go 0... yeah let's try that MAHAHAHAH!

OpenStudy (anonymous):

I guess I'm not understanding why we are searching for 0 on that portion. If A * B = 0 right, then only A OR B needs to be zero.

OpenStudy (owlcoffee):

Thing is the understanding you have for the function: \[f:f(x)= e^x\] if you look at the graph, you'll see that it has no zeroes.

OpenStudy (usukidoll):

\[0 =(2x-2x^3) (e^{(1-x^2)})\]

OpenStudy (usukidoll):

so if I substituted x = 0 \[0 =(2(0)-2(0)^3) (e^{(1-(0)^2)})\] \[0 =(2(0)-2(0)) (e^{(1-0)})\] \[0 =(0-0)) (e^{(1)})\] \[0=(0)(e)\] e is approximately 2.71 0=(0)(2.71) 0=0

OpenStudy (usukidoll):

it's like we need that \[(2x-2x^3)\] portion to help us get rid of the e ... otherwise we're stuck because x = 1 yields (2-2) -> (0)

OpenStudy (anonymous):

totally, cant escape it hey

OpenStudy (usukidoll):

I think .. splitting the problem up is only necessary to find the zeros... but we need the whole problem together and have x = 0,1,-1 in order for that equation to hold true. If we keep the split up version... something isn't going to cancel and will lead to something false.

OpenStudy (anonymous):

I think that will be an adequate solution..

OpenStudy (anonymous):

factor out the first part of the problem, and claim the 2nd term as basically irrelevant.

OpenStudy (usukidoll):

yeah.. because by setting the second equation to 0 , you will run into problems.. because ln 0 is undefined. If that's the case, just stick with the first part of the equation.

OpenStudy (usukidoll):

I've done this before when I had to find the equilibrium for Mathematical biology... if we come across something that's undefined...ignore it and focus on the problem that will give us the 0's .. in this case it was the first part of the equation which gave us x =0,1, -1.. plugging those x values back into the whole equation, will make your equation 0=0 which is true

OpenStudy (anonymous):

Sorry, I guess I should have mentioned, I just realized it might be important to know, there is a fixed domain too by the way.. it's +2 .. -2 so we dont need to worry about - infinity

OpenStudy (usukidoll):

ok no problem.. the domain lies in the x-axis

OpenStudy (usukidoll):